OpenStudy (henrietepurina):

Algebra 1 question Medal + Fan

OpenStudy (henrietepurina):

The steps to derive the quadratic formula are shown below: Step 1: ax2 + bx + c = 0 Step 2: x2 + b over ax + c over a = 0 Step 3: x2 + b over a x = - c over a Step 4: x2 +b over a x + square of quantity of b over 2 times a = -c over a + square of quantity of b over 2 times a Step 5: x plus b over 2 times a, all squared = -c over a + square of quantity of b over 2 times a Step 6: x + b over 2 times a = square root of all of negative c over a plus the square of the quantity b over 2 times a Step 7: x = - b over 2 times a ± square root of all of negative c over a plus the square of the quantity b over 4 times a = -b over 2 times a ± square root of the quantity of b squared minus 4 times a times c over 4 times a squared= negative b plus or minus square root of b squared minus 4 times a times c over 2 times a Which of the following statements best describes the first incorrect step?

OpenStudy (henrietepurina):

In Step 4, square of quantity of b over 2 times a instead of b over 2 times a was added on both sides. In Step 5, x plus b over 2 times a, all squared was written instead of x plus b over 4 times a, all squared . In Step 6, -c over a was written instead of c over a inside the root. In Step 7, b over 4 times a, all squared was written instead of square of quantity of b over 2 times a inside the root.

OpenStudy (henrietepurina):

(refresh page to remove question marks)

OpenStudy (henrietepurina):

@surjithayer

OpenStudy (henrietepurina):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

can you post a pic of the whole problem?

OpenStudy (henrietepurina):

sure

jimthompson5910 (jim_thompson5910):

I agree with your answer, it should be \(\Large \left(\frac{b}{2a}\right)^2\). Look back at step 6.

OpenStudy (henrietepurina):

exactly....

OpenStudy (henrietepurina):

so D it is?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (henrietepurina):

can you help with another @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

whats your question

OpenStudy (henrietepurina):

Which of the following tables shows the correct steps to transform x2 + 6x + 8 = 0 into the form (x - p)2 = q? [p and q are integers]

jimthompson5910 (jim_thompson5910):

you are correct again, nice work

jimthompson5910 (jim_thompson5910):

you add 1 to both sides so you have a perfect square trinomial on the left side to factor

OpenStudy (henrietepurina):

thank you :D

jimthompson5910 (jim_thompson5910):

yw